### v1

# Reframing the Relationship in Out-of-Distribution Detection

Abstract
The remarkable achievements of Large Language Models (LLMs) have captivated the attention of both academia and industry, transcending their initial role in dialogue generation. The utilization of LLMs as intermediary agents in various tasks has yielded promising results, sparking a wave of innovation in artificial intelligence. Building on these breakthroughs, we introduce a novel approach that integrates the agent paradigm into the Out-of-distribution (OOD) detection task, aiming to enhance its robustness and adaptability. Our proposed method, Concept Matching with Agent (CMA), employs neutral prompts as agents to augment the CLIP-based OOD detection process. These agents function as dynamic observers and communication hubs, interacting with both In-distribution (ID) labels and data inputs to form vector triangle relationships. This triangular framework offers a more nuanced approach than the traditional binary relationship, allowing for better separation and identification of ID and OOD inputs. Our extensive experimental results showcase the superior performance of CMA over both zero-shot and training-required methods in a diverse array of real-world scenarios.
1 Introduction
The emergence and development of Large Language Models (LLMs) [8, 84, 7, 2] have significantly reshaped the landscape of Artificial Intelligence (AI), marking a pivotal breakthrough in both academic research and practical applications. These models have not only revolutionized the way we generate conversations but also demonstrated their capacity as intermediary agents with more nuanced roles, facilitating the accomplishment of myriad tasks with unprecedented efficiency and adaptability [72, 75, 83]. The Agent paradigm has been extensively applied across multiple domains and tasks, playing a profound role [57, 58, 25].
An important and challenging task within the field of machine learning is to enhance the robustness of models across diverse scenarios. When an artificial intelligence system encounters data that significantly deviates from its training data distribution, OOD detection becomes crucial for ensuring its reliability and robustness. In the past, most OOD detection methods employed single-modal learning [79, 44, 24]. As CLIP [59] has demonstrated astonishing performance across various downstream tasks, an increasing number of CLIP-based methods for out-of-distribution (OOD) detection have emerged [61, 70, 11, 73, 22].
Question.
However, previous OOD detection methods, whether single-modal learning or CLIP-based approaches, typically rely on binary relationship to differentiate between in-distribution (ID) data and OOD data (Figure 2). These methods include solely using ID data to construct boundaries or employing a combination of ID and OOD data to demarcate their respective domains. While these methods are effective to some extent, they lack the flexibility and adaptability needed to handle the dynamic complexity of real-world data distributions.
Figure 1: Overview of Concept Matching with Agent (CMA) framework. The input image $x$ undergoes Image Encoder $\mathcal{I}$ to produce an image embedding. The concatenation of the ID labels $\mathcal{Y}_{in}$ and Agents $\mathcal{Y}_{ntc}$ is then subjected to Text Encoder $\mathcal{T}$ to generate a text embedding. The similarity between the image and text embeddings is computed, with a higher result indicating a greater degree of similarity (darker shading denotes higher similarity). This is followed by the CMA operation, which computes the $\mathcal{S}_{\text{CMA}}$ for each image as the ultimate discriminative metric. Further details are provided in Section 3.
Figure 2: Comparison between Binary Relationship and Vector Triangle Relationship in OOD detection.
Motivation.
In the realm of OOD detection, the predominant methodology consists of binary segmentation between ID and OOD data, typically facilitated by scoring functions. The effectiveness of this prototype is closely tied to the sophisticated design of these scoring functions. Drawing inspiration from the notable success of the agent-based paradigms, the comparative analysis of ID and OOD data via an agent may present a more refined and comprehensive understanding. Structurally, this paradigm shift from binary segmentation to a vector triangle relationship relational framework holds substantial potential for uncovering deeper insights, potentially transforming the interplay between ID labels and data inputs.
Our scheme.
In this paper, we propose the Concept Matching with Agent (CMA) methodology, which integrates neutral textual concept prompts in natural language as Agents within the CLIP-based OOD detection framework. Our framework is illustrated in Figure 1. These agents serve dual roles: as observers and as intermediate hubs that facilitate the interaction between ID Labels and Data inputs. By doing so, we aim to establish a vector triangle relationship among the ID labels, data inputs, and the agents themselves. In this triangular vector relationship (Figure  2), the score of OOD data is diminished due to the collision effect of the Agents, thereby widening the gap between the scores of ID data and OOD data. In other words, images closer to the ID class are less likely to be influenced by neutral text concepts, whereas images from the OOD class are more susceptible to these concepts, resulting in lower scores. The method to achieve our idea is derived from our profound insights into the Language-Vision representation (See Section 3). Building on these insights, we formulate the entire Concept Matching with Agent (CMA) framework.
In summary, our proposed method, CMA, possesses three distinct advantages. (1) Our approach obviates the need for training data, enabling zero-shot OOD detection while leveraging the collision between Agent and both ID and OOD data to effectively widen the gap between the two, ensuring robust performance. This stands in stark contrast to traditional OOD detection methods, which rely on extensive external data for intricate training [79, 44]. (2) Our method exhibits remarkable scalability, allowing for the tailoring of specialized Agents to suit various scenarios, thereby further enhancing performance (See Section 5). This is facilitated by the flexibility of our triangular vector relationship, which can enhance the impact of certain OOD images through specific Agents, thereby reducing their scores. (3) It is noteworthy that the CMA maintains robustness against both hard OOD inputs, encompassing both semantic hard OODs [74] and spurious OODs [50]. This makes our approach a truly practical and viable option. The contributions of our study are summarized as follows:
  • Drawing upon the concept of Agents in LLMs, we propose the incorporation of agent-based observation into OOD detection. By facilitating interactions among agents, ID labels, and data inputs, we establish a vector triangle relationship for them. This structured shift diverges from the conventional binary frameworks, offering enhanced flexibility in application scenarios and providing a novel analytical perspective on OOD detection.
  • We propose a novel CLIP-based OOD Detection framework. Compared to previous methods, our approach more effectively achieves the objective of OOD Detection: it widens the gap between ID and OOD, and possesses enhanced versatility and practicality.
  • We conducted experiments on various datasets with distinct ID scenarios and demonstrated that CMA achieves superior performance across a wide range of real-world tasks. Compared to most existing OOD detection methods, CMA brings substantial improvements to the large-scale ImageNet OOD benchmark.
2 Preliminaries
Contrastive Vision-Language Models
In comparison to traditional CNN architectures, the ViT [16] leverages the Transformer Encoder framework [68] to accomplish the task of image classification, realizing the possibility of utilizing language model architectures for visual tasks. This provides insights into the field of vision-language representation learning, with CLIP [59] being a notable representative. CLIP employs self-supervised contrastive objectives to embed images and their corresponding textual descriptions into a shared feature space, achieving alignment between the two. Structurally, CLIP, which utilizes a dual-stream architecture, comprises an image encoder $\mathcal{I}:x\to\mathbb{R}^{d}$ and a text encoder $\mathcal{T}:t\to\mathbb{R}^{d}$. After pretraining on a dataset of 400 million text-image pairs, the joint visual-language embedding of CLIP associates objects in various patterns. Due to the robust performance of CLIP, several OOD detection methods based on it have emerged. Nevertheless, the challenge of how to better utilize Language-Vision representation for OOD detection remains a difficult yet significant issue.
Zero-shot OOD Detection
For traditional OOD detection frameworks [24], a common assumption is made of a typical real-world scenario wherein classifier $f$ are trained on ID data categorized as $\mathcal{Y}_{in}=\left\{1,...,C\right\}$ and subsequently deployed in an environment containing samples from unknown classes $y\notin\mathcal{Y}_{in}$, outside the distribution of the ID. The classifier $f$ is then tasked with determining membership to the ID. In contrast, zero-shot OOD detection [48, 51, 21], in line with the current trend of deep learning, leverages pre-trained models on open datasets, eliminating the need for additional training of the model. The approach determines membership to the ID by calculating results through mapping the data into a common space $\mathbb{R}^{d}$. Compared to traditional OOD detection frameworks that necessitate training, the zero-shot OOD detection framework is notably more versatile and practical.
3 Method
3.1 CLIP-based OOD Detection
The CLIP model aligns image features with text features describing the image in a high-dimensional space by simultaneously training image and text encoders on a large dataset, thereby learning rich visual-language joint representations. When applied to OOD detection tasks, CLIP only requires class names and does not require training on specific ID data, allowing it to attempt to classify or determine whether an input image belongs to a known class. It is worth noting that the ID classes in CLIP-based OOD detection refer to the classes used in downstream classification tasks, which are different from the pre-trained classes in the upstream. The OOD classes are those that do not belong to any of the ID classes in the downstream tasks.
For current CLIP-based OOD detection, MCM [48] has become a basic paradigm. Its core idea is to treat text embeddings as "concept prototypes" and evaluate their in-distribution or out-of-distribution properties by measuring the similarity between the input image features and these concept prototypes. Specifically, given a set of ID categories $\mathcal{Y}_{in}$ with a corresponding text description for each category, we first use a pre-trained text encoder $\mathcal{T}$ to convert these text descriptions into $d$ dimensional vectors $\mathbf{c}_{i}=\mathcal{T}(t_{i})\in\mathbb{R}^{d}$, where $i\in\{1,2,...,N\}$ and $N$ is the number of ID categories. For any input image $\mathbf{x}^{\prime}$ whose visual features are extracted by an image encoder $\mathcal{I}$ as $\mathbf{v}^{\prime}=\mathcal{I}(\mathbf{x}^{\prime})\in\mathbb{R}^{d}$, the MCM score is defined as the cosine similarity between the visual features and the closest concept prototype, which is scaled by an appropriate softmax to enhance the separability of ID and OOD samples:
$\mathcal{S}_{\text{MCM}}(\mathbf{x}^{\prime};\mathcal{Y}_{in},\mathcal{T},\mathcal{I})=\frac{\exp(\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{\hat{y}})/\tau)}{\sum_{i=1}^{N}\exp(\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})/\tau)},$
where $\text{sim}(\cdot,\cdot)$ represents cosine similarity, $\hat{y}=\arg\max_{i}\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})$ represents the most matching concept index, and $\tau$ is a temperature parameter used to adjust the distribution of similarity.
3.2 How to construct vector triangle relationships using Agents
Figure 3: Heatmaps depicting the cosine similarity between image inputs and ID concept vectors. In Figure (A), the ID concept vectors consist of sentences containing the word "cat" of varying lengths. It is observed that images tend to align with longer sentences regardless of whether there is a matching ID concept. Concurrently, keywords such as "white" significantly influence image matching. In Figure (B), aside from "cat", no ID concept can be precisely matched with the given images. However, other than cat images, all images exhibit a preference for aligning with a long sentence devoid of tangible objects. Notably, cat image remains unaffected, aligning solely with the ID concept "cat".
Although the employment of the Agent paradigm in the realm pertaining to LLMs has become a matter of course, the challenge lies in its adaptation to the domain of Out-of-distribution Detection. To address this issue, we have conducted a systematic examination of linguistic visual representations and, based on this, conducted targeted experiments, identifying three primary phenomena.Figure  3 shows two basic examples. These laws provide a new perspective for us to understand and optimize multimodal learning models:
  • The length of prompt will affect the prediction. In similar text descriptions, the longer the prompt is within a certain range, the higher the score.
  • Different words have different weights in the text description. The more important the words are to the overall description of the picture, the higher the weight.
  • Neutral prompts have little impact on images in the ID category. Conversely, for images in the OOD category, neutral prompts can significantly reduce the score of the ID textual description.
The causes of these phenomena can be contemplated through the training methods and mechanisms of the CLIP-like models.Given that the common text input for pre-trained models during training is long sentences, they possess higher confidence in inferring long sentences. Additionally, CLIP-liked models undergo contrastive learning training, which enables the model to autonomously focus on the differences between various text descriptions, thereby giving higher weights to key words and skewing the predicted outcomes towards the positive class without being influenced by other words.
Drawing on these significant observations, we propose the inclusion of neutral prompts unrelated to the ID category as Agents. By engaging Agents in a ’collision’ with OOD and ID data, we aim to distance the OOD data from the ID domain.
3.3 Proposed approach
Based on the aforementioned analysis, Concept Matching with Agent (CMA) employs neutral prompts as Agents to widen the gap between ID and OOD by constructing a vector triangle relationship. Essentially, CMA employs external agents to minimize the maximum score, which is fundamentally distinct from traditional learning methods. Therefore, it does not require additional data or training. Specifically, for a set of textual descriptions $t_{i}$, where $t\in\{1,2,...,N\}$ corresponding to ID categories $\mathcal{Y}_{in}$, we select a certain number of neutral prompts $\mathcal{Y}_{ntc}$ to concatenate with them. The number is generally the same as the number of ID categories $N$, and the concatenated textual concepts $\mathcal{Y}_{con}$ are used as the input of the text encoder $\mathcal{T}$ to output text features $\mathbf{c}_{i}=\mathcal{T}(t_{i})$, where $i\in\{1,2,...,2N\}$, $t_{i}\in\mathcal{Y}_{con}=\{\mathcal{Y}_{in},\mathcal{Y}_{ntc}\}$. Then, it is calculated with the image features output $\mathbf{v}^{\prime}=\mathcal{I}(\mathbf{x}^{\prime})$ by the image encoder $\mathcal{I}$.
Formally, we define the CMA score as:
$\mathcal{S}_{\text{CMA}}(\mathbf{x}^{\prime};\mathcal{Y}_{in},\mathcal{Y}_{ntc},\mathcal{T},\mathcal{I})=\frac{\exp(\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{\hat{y}})/\tau)}{\sum_{i=1}^{2N}\exp(\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})/\tau)},$
where $\text{sim}(\cdot,\cdot)$ represents cosine similarity, $\hat{y}=\mathop{\arg\max}\limits_{1\leq i\leq N}\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})$ represents the most matching concept index in ID categories, and $\tau$ is a temperature parameter used to adjust the distribution of similarity.
It is particularly noteworthy that although our calculation method is similar to MCM, there are significant differences in detail: (1) When calculating the highest score, the scores corresponding to neutral prompts should be excluded from the calculation, but when performing softmax scaling, the scores corresponding to neutral prompts should be retained; (2) For text descriptions of ID categories, we do not apply prompts for text enhancement, but only use the corresponding category names. This is to further differentiate between ID and neutral prompts, making the model more focused on whether the image truly corresponds to the ID category.
Finally, our OOD detection function can be formally formulated as:
$g\left(\mathbf{x}^{\prime};\mathcal{Y}_{\mathrm{in}},\mathcal{Y}_{ntc},\mathcal{T},\mathcal{I}\right)=\left\{\begin{array}[]{ll}1,&S_{\mathrm{CMA}}\left(\mathbf{x}^{\prime};\mathcal{Y}_{\text{in }},\mathcal{Y}_{ntc},\mathcal{T},\mathcal{I}\right)\geq\lambda\\ 0,&otherwise\end{array},\right.$
where 1 indicates ID and 0 indicates OOD. $\lambda$ is the threshold, and examples below $\lambda$ are considered OOD inputs.
4 Experiments
Table 1: Comparison results on ImageNet-1k OOD benchmarks. We use ImageNet-1k as ID dataset. All methods use CLIP-B/16 as a backbone. Bold values represent the highest performance.
iNaturalist SUN Places Texture Average
[…]
Requires training (or w. fine-tuning)
MSP [24] 40.89 88.63 65.81 81.24 67.90 80.14 64.96 78.16 59.89 82.04
Energy [45] 21.59 95.99 34.28 93.15 36.64 91.82 51.18 88.09 35.92 92.26
ODIN [42] 30.22 94.65 54.04 87.17 55.06 85.54 51.67 87.85 47.75 88.80
ViM [71] 32.19 93.16 54.01 87.19 60.67 83.75 53.94 87.18 50.20 87.82
KNN [65] 29.17 94.52 35.62 92.67 39.61 91.02 64.35 85.67 42.19 90.97
NPOS [66] 16.58 96.19 43.77 90.44 45.27 89.44 46.12 88.80 37.93 91.22
CoOp [86] 43.38 91.26 38.53 91.95 46.68 89.09 50.64 87.83 44.81 90.03
LoCoOp [52] 38.49 92.49 33.27 93.67 39.23 91.07 49.25 89.13 40.17 91.53
Zero-shot (no training required)
MCM [48] 30.94 94.61 37.67 92.56 44.76 89.76 57.91 86.10 42.82 90.76
GL-MCM [51] 15.18 96.71 30.42 93.09 38.85 89.90 57.93 83.63 35.47 90.83
CMA(Ours) 23.84 96.89 30.11 93.69 29.86 93.17 47.35 88.47 32.79 93.05
Table 2: Zero-shot OOD detection with CMA based on CLIP-B/16 with various ID datasets.
iNaturalist SUN Places Texture Average
[…]
FashionMNIST [76] 0.00 100.00 0.00 100.00 0.00 100.00 13.61 93.60 3.40 98.40
STL10 [10] 0.00 100.00 2.56 99.01 0.00 100.00 0.00 100.00 0.64 99.75
OxfordIIIPet [55] 0.00 99.89 0.37 99.89 0.96 99.71 0.35 99.85 0.42 99.84
Food101 [6] 0.25 99.90 0.49 99.86 1.79 99.40 2.99 99.33 1.38 99.62
CUB-200 [69] 0.00 100.00 0.00 100.00 0.00 99.99 0.00 100.00 0.00 100.00
PlantVillage [30] 2.95 97.83 0.18 98.11 1.12 98.41 4.37 97.56 2.16 97.98
LFW [27] 2.89 99.14 2.24 99.52 8.04 98.25 18.88 95.34 8.01 98.06
Stanford-dogs [34] 0.11 99.92 0.16 99.91 0.58 99.73 0.67 99.72 0.38 99.82
FGVC-Aircraft [47] 0.00 99.99 0.67 99.87 1.02 99.69 0.00 99.99 0.42 99.89
Grocery Store [36] 0.03 99.89 0.15 99.97 0.69 99.82 0.79 99.76 0.42 99.86
CIFAR10 [37] 0.00 100.00 5.12 98.29 2.56 99.67 0.00 99.88 1.92 99.46
CIFAR100 [37] 12.80 97.15 17.92 95.66 15.36 96.19 4.53 98.42 12.65 96.86
Table 3: Comparison results on hard OOD detection tasks.
Method ID ImageNet-10 ImageNet-20 ImageNet-10 ImageNet-20 Waterbirds
OOD ImageNet-20 ImageNet-10 ImageNet-100 ImageNet-100 Spurious OOD
FPR95 / AUROC FPR95 / AUROC FPR95 / AUROC FPR95 / AUROC FPR95 / AUROC
MSP [24] 9.38 / 98.31 12.51 / 97.70 5.82 / 98.88 14.14 / 95.73 39.57 / 90.99
MCM [48] 5.00 / 98.71 12.91 / 98.09 3.70 / 99.09 13.16 / 96.32 5.87 / 98.36
CMA(Ours) 3.1 / 99.19 6.2 / 98.71 2.94 / 99.18 10.32 / 96.64 3.22 / 99.01
4.1 Setup
Datasets
We conducted a comprehensive evaluation of the performance of our method across various dimensions and compared it with widely employed OOD detection algorithms. (1) We assessed our approach on the ImageNet-1k OOD benchmark. This benchmark utilizes the large-scale visual dataset ImageNet-1k [13] as the ID data and four OOD datasets (including subsets of iNaturalist [67], SUN [77], Places[85], and Textures [9], which are same as Sun et al. [65]) to fully evaluate the method’s performance across various semantic and scenario contexts. (2) We evaluated our method on various small-scale datasets. Specifically, we considered the following ID datasets: FashionMNIST [76], STL10 [10], OxfordIIIPet [55], Food-101 [6], CUB-200 [69], PlantVillage [30], LFW [27], Stanford-dogs [34], FGVC-Aircraft [47], Grocery Store [36], and CIFAR-10 [37]. (3) We assessed our method on hard OOD tasks [74, 50]. Following the standards of the MCM [48], we evaluated using subsets of ImageNet-1k, namely ImageNet-10 and ImageNet-20, which have similar classes (e.g., dog (ID) vs. wolf (OOD)). During the experiments, we ensured that each OOD dataset did not overlap with the ID dataset in terms of classes.
Model
In our experiment, all algorithms uniformly employ CLIP [59] as the pre-trained model, which is one of the most prevalent and publicly available visual-linguistic models. Specifically, we utilize CLIP-B/16 as the foundational evaluation model, consisting of a ViT-B/16 transformer [16] serving as the image encoder and a masked self-attention Transformer [68] as the text encoder. Additionally, unless otherwise specified, the temperature coefficient is uniformly set to 1 across all algorithms.
Metrics
For evaluation, we use the following metrics: (1) the false positive rate (FPR95) of OOD samples when the true positive rate of in-distribution samples is at 95%, (2) the area under the receiver operating characteristic curve (AUROC). All evaluation outcomes for our method are derived from the average of three experiments, and (3) ID classification accuracy (ID ACC).
4.2 Main Results
OOD detection on Large-scale datasets.
The benchmarking of large-scale OOD datasets demonstrates the feasibility of the method for real-world applications and holds significant value. Typically, we employ ImageNet-1k as the ID dataset for Large-scale OOD detection. Table  3 presents the performance of our method in comparison with other approaches under this benchmark. Overall, our method surpasses other methods, achieving superior performance. When juxtaposed against the average performance of Zero-shot methods, CMA demonstrates enhancements of 2.22% in terms of AUROC and 2.68% in terms of FPR95. Similarly, when juxtaposed against the average performance of Require training methods, CMA also demonstrates improvements, achieving enhancements of 0.79% in AUROC and 3.13% in FPR95, thereby showcasing its exceptional performance.
OOD detection on small-scale datasets.
In contrast to benchmarks based on large-scale OOD datasets, small-scale OOD detection often features fewer distinct ID categories. Demonstrating robust performance across these categories is indicative of a method’s scalability. Table  3 illustrates the efficacy of our approach across various ID datasets. A notable outcome is that our method achieves impressive performance across these datasets, especially when no specific training was tailored to each dataset.
Hard OOD detection.
Table  3 presents the outcomes of various zero-shot approaches in hard out-of-distribution (OOD) detection. For dataset configurations, we alternated between Imagenet-10 and Imagenet-20 (See Appendix A), which share similar semantic content, as the ID and OOD datasets, respectively. Alternatively, we employed the setup of using Imagenet-10 and Imagenet-20 as the ID datasets, with Imagenet-100 serving as the OOD dataset. Additionally, to account for spurious correlations, we utilized the Spurious OOD detection benchmark introduced by Ming et al. [48] and conducted comparative experiments on the Waterbirds dataset [50]. Across various experiments, CMA outperformed MSP and MCM in terms of AUROC and FPR95.
4.3 Ablation study
Whether to use softmax scaling.
Table 4: Comparison results of CMA with and without softmax scaling based on CLIP-B/16. The ID dataset is ImageNet-1k.
iNaturalist SUN Places Texture Average
[…]
w/o softmax scaling 70.27 85.64 57.75 88.49 46.31 89.80 80.67 77.75 63.75 85.42
w/ softmax scaling 23.84 96.89 30.11 93.69 29.86 93.17 47.35 88.47 32.79 93.05
Table 5: Comparison results of CMA with different models. The ID dataset is ImageNet-1k.
iNaturalist SUN Places Texture Average
[…]
CLIP-RN50 24.08 95.25 37.54 91.32 44.53 88.92 39.21 90.84 36.34 91.58
CLIP-B/16 23.84 96.89 30.11 93.69 29.86 93.17 47.35 88.47 32.79 93.05
CLIP-L/14 19.52 96.34 25.78 94.69 26.54 94.29 49.48 88.21 30.33 93.38
Although MCM has demonstrated the significant role of softmax scaling in CLIP-based OOD detection, our method CMA underscores the even more crucial role of softmax scaling. Table  5 illustrates the disparities between the use and non-use of softmax scaling. A crucial factor is that without softmax scaling, the score formula in CMA would become $\mathcal{S}_{\text{CMA}}^{w/o}=\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{\hat{y}})/\tau$, where $\hat{y}=\mathop{\arg\max}\limits_{1\leq i\leq N}\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})$ represents the most matching concept index in ID categories. In comparison to the original formula, the absence of softmax scaling essentially excludes Agents from the final score calculation, resulting in a score of $\mathcal{S}_{\text{CMA}}^{w/o}=\max_{i\in[N]}\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})/\tau$. This prevents the construction of vector triangle relationships, thus negating the impact of Agents and leading to a significant reduction in performance. This also validates the rationality and effectiveness of $\mathcal{S}_{\text{CMA}}$.
Comparative performance across models.
We conducted a comparative analysis of the performance of CMA across various models [59], with the results presented in Table  5. The findings indicate that as the backbone model’s performance improves, the effectiveness of CMA also intensifies, resulting in superior OOD detection capabilities. Furthermore, this study underscores the fact that our approach is not reliant on a single model, but can be effectively integrated across diverse models, demonstrating remarkable scalability.
5 Discussion
Figure 4: Impact curve of $k$ on the performance of CMA. $k=\frac{\text{number of Agents}}{\text{number of ID Labels}}$
Figure 5: Comparison with different Agents.
The number of Agents.
In section 3, we present the score calculation formula $\mathcal{S}_{\text{CMA}}(\mathbf{x}^{\prime};\mathcal{Y}_{in},\mathcal{Y}_{ntc},\mathcal{T},\mathcal{I})$ for CMA, where we set the number of Agents to $N$, aligning with the number of ID categories. This number was determined based on extensive experiments, with $k$ set to $k=\frac{M}{N}$, where $M$ is the number of Agents.Concurrently, the formula for calculating the score has been modified to $\mathcal{S}_{\text{CMA}}^{\prime}(\mathbf{x}^{\prime};\mathcal{Y}_{in},\mathcal{Y}_{ntc},\mathcal{T},\mathcal{I})=\frac{\exp(\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{\hat{y}})/\tau)}{\sum_{i=1}^{N}\exp(\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})/\tau)+\sum_{i=1}^{M}\exp(\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})/\tau)}$, where $\hat{y}=\mathop{\arg\max}\limits_{1\leq i\leq N}\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})$. Experiments were conducted for various values of $k$. Figure 4 illustrates the results of our experiments conducted on ImageNet-1k. The optimal value of $k$ was identified as 1. Additionally, it was observed that an excessive number of Agents does not enhance the overall performance of CMA, whereas a insufficient number of Agents can adversely affect it. This finding illuminates the rationality of the vector triangle relationship. An insufficient number of Agents fails to exert the same impact on OOD data as it does on ID categories, leading to structural disarray and performance degradation. Once the number of Agents reaches a certain threshold, a stable structure is established, after which increasing the number of Agents does not significantly improve performance.
Performance of different Agents.
Figure 5 illustrates the impact of different neutral terms on the OOD detection performance of agents. It is evident that the impact of different agents on performance is substantial. On average, the best and worst outcomes in terms of FPR95 and AUROC can differ by 19.95% and 4.50%, respectively. From another perspective, certain agents excel in specific contexts, such as "a large photo of the scene we live in." While their overall performance across multiple environments lags significantly behind the optimal level, they achieve their best results when using the SUN dataset [77] as OOD data. On one hand, this highlights a potential limitation of CMA, where different agents can yield disparate outcomes. On the other hand, it underscores the significant potential of CMA, as it can tailor specific agents to suit diverse environments, thereby enhancing performance.
6 Related Works
Single-modal Out-of-distribution detection
Visual Out-of-distribution (OOD) detection [79, 44, 80], a significant area in machine learning systems and computer vision, involves discerning images devoid of specific semantic information. Following a prolonged period of development, this field has evolved into two broad categories: single-modality OOD detection methods and multimodal OOD detection approaches that leverage semantic information. For the former, D Hendrycks et al. [24] outlines the fundamental paradigm within neural networks, which has been the foundation for numerous subsequent developments, including output-based methods [4, 15, 17, 23, 26, 28, 45, 64], density-based methods [1, 88, 12, 63, 56], distance-based methods [38, 49, 62, 65], reconstruction-based methods [39, 14, 87, 82], and gradient-based methods [42, 29, 31]. Some works [53, 20, 5] have also provided enhanced theoretical analyses for OOD detection. Recent studies have sought to reexamine and address the OOD detection problem from various perspectives. Ammar et al. [3] proposes utilizing "neural collapse" and the geometric properties of principal component spaces to identify OOD data. Lu et al. [46] models each class with multiple prototypes to learn more compact sample embeddings, thereby enhancing OOD detection capabilities.
CLIP-based Out-of-distribution detection
As multimodal models emerge as a new trend, multimodal pretrained models [59, 32, 78] are increasingly being introduced into other domains [81, 40, 35, 60, 43, 19]. The idea of utilizing large-scale pretrained multimodal models that rely solely on outlier class names without any accompanying images was first proposed in Fort et al. [21]. ZOC [18] employs an extended model to generate candidate unknown class names for each test sample. MCM[48], for the first time, abandons the traditional reliance on sample assumptions, starting from a language-vision representation and using a temperature scaling strategy and the maximum predicted softmax value as the out-of-distribution (OOD) score, establishing a paradigm for CLIP-based OOD detection. Building on MCM, GL-MCM [51] enhances the model by aligning the global and local visual-textual features of CLIP [59]. LoCoOp[52] proposes a few-shot OOD detection method within the framework of prompt learning. LSN [54], NegPrompt [41], and NegLabel [33] all enhance the effectiveness of OOD detection by introducing negative prompts and utilizing methods such as prompt learning and sampling. Our approach, CMA, is rooted in MCM and offers a fresh perspective on CLIP-based OOD detection from the language-vision representation standpoint. It is a post-hoc detection method that obviates the need for learning and external data.
7 Conclusion
This paper proposes a novel zero-shot OOD detection framework, Concept Matching with Agent (CMA). By introducing the concept of Agents into the OOD detection task, a vector triangular relationship consisting of ID labels, data inputs, and Agents is constructed, offering a fresh perspective on OOD detection. Beginning with a language-vision representation, we demonstrate the impact of neutral words on CLIP-liked models, thereby proposing the innovative idea of treating neutral words as Agents. Building on this, we propose a novel score computation method based on CMA. By incorporating the relationships between Agents and data inputs, this method enables unique interactions between ID data and OOD data with Agents, thereby facilitating a better separation between ID and OOD. We investigate the effectiveness of CMA across various scenarios, including large-scale datasets, small-scale datasets, and hard OOD detection, achieving outstanding performance across a wide range of tasks. Finally, we delve deeper into CMA, highlighting its flexibility and scalability. We hope that our work will inspire future exploration of new paradigms for OOD detection.
Appendix A Experimental Details
A.1 Software and Hardware
All methods are implemented in Pytorch 2.0.1 and Python 3.8. We run all OOD detection experiments on 4 NVIDIA RTX A6000 GPUs with 10 Intel Xeon Gold 5220 CPUs.
A.2 Datasets
ImageNet-10 We use the ImageNet-10 in [48] that mimics the class distribution of CIFAR-10 but with high-resolution images. It contains the following categories (with class ID): warplane (n04552348), sports car (n04285008), brambling bird, (n01530575), Siamese cat (n02123597), antelope (n02422699), Swiss mountain dog (n02107574), bull frog (n01641577), garbage truck (n03417042), horse (n02389026), container ship (n03095699).
ImageNet-20 We use the ImageNet-20 in [48], which consists of 20 classes semantically similar to ImageNet-10 (e.g. dog (ID) vs. wolf (OOD)). It contains the following categories: sailboat (n04147183), canoe (n02951358), balloon (n02782093), tank (n04389033), missile (n03773504), bullet train (n02917067), starfish (n02317335), spotted salamander (n01632458), common newt (n01630670), zebra (n01631663), frilled lizard (n02391049), green lizard (n01693334), African crocodile (n01697457), Arctic fox (n02120079), timber wolf (n02114367), brown bear (n02132136), moped (n03785016), steam locomotive (n04310018), space shuttle (n04266014), snowmobile (n04252077).
ImageNet-100 We use the ImageNet-100 in [48], which is curated from ImageNet-1k.
The aforementioned datasets can be accessed at https://github.com/deeplearning-wisc/MCM.
A.3 Sources of model checkpoints
For CLIP models, our reported results are based on checkpoints provided by Hugging Face for CLIP-B/16 https://huggingface.co/openai/clip-vit-base-patch16 and CLIP-L/14 https://huggingface.co/openai/clip-vit-large-patch14. Similar results can be obtained with checkpoints in the codebase by OpenAI https://github.com/openai/CLIP. Note that for CLIP-RN50, which is not available in Hugging Face, we use the checkpoint provided by OpenAI.
A.4 Hyperparameters
In our approach, in addition to the number of agents, $M$, which we have already discussed in the Discussion section (See Section 5), an important hyperparameter is temperature parameter $\tau$. In all the experiments presented in the main text, $\tau$ is set to 1. This is because our empirical findings (Table 6) indicate that CMA is insensitive to $\tau$, resulting in minimal variations in outcomes.
Appendix B ID Classification Accuracy
Table 7 presents the multi-class classification accuracies achieved by various methods on ImageNet-1k. The findings indicate that CMA does not compromise the model’s accuracy on the original classification task.